OpenStudy (anonymous):

Find the first and second derivative of the function. Check to see that your answers are reasonable by comparing the graphs of f, f ', and f ''. f(x) = 3x − 4x^7/8

OpenStudy (anonymous):

f'(x)=\[3-\frac{ 7 }{ 2 }x ^{-\frac{ 7 }{ 8 }}\]

OpenStudy (campbell_st):

not quite the power of the derivative is (n - 1) in your question its (7/8 - 1)

OpenStudy (anonymous):

-1/8

OpenStudy (anonymous):

f''(x)= \[\frac{ 7 }{ 16 }x ^{-\frac{ 9 }{ 8 }}\]

OpenStudy (campbell_st):

looks ok... well done